### v1

# Toward Errorless Training ImageNet-1k

Abstract: In this paper, we describe a feedforward artificial neural network trained on the ImageNet 2012 contest dataset [7] with the new method of [5] to an accuracy rate of 98.3% with a 99.69 Top-1 rate, and an average of 285.9 labels that are perfectly classified over the 10 batch partitions of the dataset. The best performing model uses 322,430,160 parameters, with 4 decimal places precision. We conjecture that the reason our model does not achieve a 100% accuracy rate is due to a double-labeling problem, by which there are duplicate images in the dataset with different labels.
1. Introduction
Artificial intelligence (AI), and intelligence in general, should exhibit these traits at a minimum: predictability, creativity, and learning. For AI models built on neural networks, the predictability is based on the continuity of all transformations in the networks as building blocks. Namely, every parameter and variable has its neighborhood of continuity in which the model behaves similarly, leading to the theoretical basis for predictability. All large language models are capable of hallucinating, and hence are creative by definition. The third capability, however, can be defined rigorously. By definition, learning is not making the same mistake twice. In practice, this means that a model can be trained error-free every time. This criterion is both necessary and sufficient. In fact, for discrete functions, errorless training has been guaranteed theoretically by the discrete version of the Universal Approximation Theorem (UAT, Theorem 2.5 of [9]) since 1989.
However, no methods for errorless training were known til 2023, when the first author presented a new algorithm, called gradient descent tunneling (GDT) in [5]. The method uses any stochastic gradient descent (SGD) algorithm to train a feedforward neural network (FNN, [16, 1]) to a modest accuracy and then follows it up by training the model to zero error rate one data batch at a time, using the idea of homotopy ([2]) from Numerical Analysis and Dynamical Systems, fields far from Machine Learning (ML). To demonstrate the effectiveness of GDT, [5] trained several error-free models on the MNIST benchmark data for classifying handwritten digits [13, 14]. The model weights for error-free FNNs are provided in the GitHub repository [6] for the MNIST, as well as for the FashionMNIST ([14, 19]) by the second author. These datasets are relatively small in size. The question is, can GDT be effective for large data sets? An affirmative answer to this question has practical import for predictive precision AI.
The purpose of this paper is to report the results of applying the GDT method to the ImageNet-1k dataset [7]. We will present a simple model architecture with the aim of training the dataset error-free or, at the very least, achieving high accuracy rates. The model architecture is similar to a mixture of experts model [18], consisting of multiple FNN feature modules in parallel. We will present evaluation results on the model’s performance, which is independent of training, and evaluations on the model’s training, which only requires the model’s weight and bias parameters. The highest performing model has 98.3% accuracy with a 99.969% Top-1 accuracy rate, and among its constituent modules, the lowest accuracy rate is 98.965% and the highest accuracy rate is 100%. There are 285.9 labels on average for each of the 10 batches of the dataset that achieve the 100% accuracy rate. The best performing model uses 322,430,160 parameters and 1,606,160 neurons. Our results also suggest that the reason for our model’s imperfect performance is due to possible labeling inconsistencies in the ImageNet-1k dataset. Specifically, images may appear multiple times in the dataset with different labels.
Before our model, according to the benchmark tracking website Papers with Code, [21] developed the most accurate model trained on the ImageNet-1k dataset with $91.0\%$ accuracy and 2.1 billion parameters. The model with the fewest parameters was [8], which attained $65.59\%$ accuracy and 1.03 million parameters. See also [12] for the AlexNet, which achieved a Top-1 37.5% error rate with 60 million parameters and 650,000 neurons.
With the utilization of GPUs, many advancements in model architecture have been made that greatly increase the accuracy of machine learning models, e.g., [15, 3, 11], and many others. Still, the approach has largely remained the same: do some versions of stochastic gradient descent until the model reaches a sufficiently accurate benchmark ([17, 13, 14, 10]). We believe that our approach presented in this report should provide us with a novel training alternative in the fields of ML and AI.
2. Model Architecture
A model of image classification, $M$, is a mapping that takes in an image as input and outputs a label. Our model is structured as follows. First, a fixed number of features of the image are selected for a model. For example, for a color image, a feature can be one of the three primary color channels, red (R), green (G), and blue (B), or a combination of the RGB channels, such as the Y channel $Y=0.2126R+0.7152G+0.0722B$. A combination of selected $n$ features uniquely defines a model, referred to as a featured model. Seventeen different features are listed in Tab.1. The set of the RG weights for RGg1, 0.618 and 0.382, comes from the Golden ratio. This choice was based on observing that the primary color features are generally easier to train than the features with equal weights – e.g., RG and eRGB. RGg2 is the symmetric feature of RGg1. Feature BW is a popular choice for the luminance of color images. The RGB color space conversion to the XYZ color space is given by the last $3\times 3$ matrix.
feature RGB Weights
R 1 0 0
G 0 1 0
B 0 0 1
RGg1 0.618 0.382 0
RBg1 0.618 0 0.382
GBg1 0 0.618 0.382
RGg2 0.382 0.618 0
RBg2 0.382 0 0.618
GBg2 0 0.382 0.618
RG 0.5 0.5 0
RB 0.5 0 0.5
GB 0 0.5 0.5
eRGB 1/3 1/3 1/3
BW 0.299 0.587 0.114
X 0.4125 0.3576 0.1804
Y 0.2126 0.7152 0.0722
Z 0.0193 0.1192 0.9502
Table 1. RGB Features
Next, we subject every feature of the model to a determination of labeling. To do so, we partition the set of labels into $k$ groups, which are referred to as modules. For the ImageNek-1k dataset, we partition the 1000 labels into $k=40$ (primary) modules, each containing 25 labels. For example, module one ($m_{1}$) consists of labels from 0 to 24, and $m_{40}$ consists of labels from 960 to 999.
For each module $m_{j}$, we train $r$ FNNs, $n_{j1},\dots,n_{jr}$, see Fig.1 for an illustration of the architecture. For the first module $m_{1}$ of labels numerated from 0 to 24, the output of $m_{1}$’s $r$ FNNs, $n_{1s}$, is a label between 0 and 24. We will use the cross-entropy for the loss function and the nearest-neighbor protocol for classification.
As a result, for every input image, each FNN will give a prediction in its module labels with the minimum cross-entropy loss. In total, there are $n\times k\times r$ predicted labels, together with their nearest-neighbor cross-entropy losses. We will denote the labels and the losses by $\ell_{i,j,s},\ \varepsilon_{i,j,s}]$ with $1\leq i\leq n$ for the feature number, $1\leq j\leq k$ for the module number, and $1\leq s\leq r$ for the FNN number.
The last operation of the model applies the majority voting method to determine a final prediction from the collection of predicted labels $\{\ell_{i,j,s}\}$ by finding the minimum of the losses $\{\varepsilon_{i,j,s}\}$. In the case of a tie, we use the smallest variance of the losses to break the tie. Because the values of the losses are real numbers with many decimal places, the smallest variance is almost surely unique.
Figure 1. Model Architecture. Every branch from the input to the output is an FNN. Each branch is uniquely defined by the triplet $(i,j,k)$ for feature $f_{i}$, module $m_{j}$, and sub-module $n_{js}$. The corresponding FNN is $N_{i,j,s}$ with transformed training batch $B_{i,j,s}$. The model is a parallel collection of $n\times k\times r$ FNNs.
To complete the description of the model architecture, we use a standard FNN with one or two hidden layers for all $n_{js}$ of all features. The number of nodes for the output layer is $1000/k$, which equals 25 in our case since we take $k=40$. The first hidden layer has 256 nodes, and the second, if present, has 77 nodes. The number of nodes for the input layer is 900, which we explain in §3. As a result, we use $900\times 256\times 25$ architecture for FNNs with one hidden layer, and $900\times 256\times 77\times 25$ for FNNs with two hidden layers.
Notation-wise, we will use S_hxx_myy for SGD trained models and T_hxx_myy for GDT trained models, with all 17 RGB features described in Tab.1. To distinguish the usage of featured models later, we will call them proto-model. Here, $x$ with $x=1,2$ is for the number of hidden layer(s), and $y$ with $y=1,2$ is for different proto-models. That is, the first capital letters, S or T, are for the type of training, and the integer index $y$ is for distinct proto-models trained by the methodology S or T, with the hidden architecture hxx. For example, S_h2_m1 is the first SGD proto-model with 2 hidden layers. With one hidden layer, each proto-model with $n=17,k=40,r=2$ has $17\times 40\times 2\times(900+256+25)=$ 1,606,160 neurons and $17\times 40\times 2\times(901\times 256+257\times 25)=$ 322,430,160 many parameters for edge weight matrices $W$ of sizes $900\times 256$ and $256\times 25$ and bias vectors $b$ of sizes $256\times 1$ and $25\times 1$ for the hidden layer and the output layer respectively.
3. Training
Training is carried out for the proto-models. There are more than 1.2 million images in the ImageNet-1k dataset, with 1000 labels. To train all FNNs, we first organize the images into $k$ batches, with $k$ the same as the number of modules. Images from a batch have the same labels as those that are used to train the corresponding module. For example, batch 1 contains all images with labels from 0 to 24, and batch 40 contains all images with labels from 960 to 999. On average, each batch contains about 30,000 images. The reason for having multiple FNNs for each module/batch is for easier training of the FNNs. In our case, we use $r=2$ FNNs for each module/batch so that FNN $n_{j1}$ is trained on the first half images of the batch $m_{j}$, and FNN $n_{j2}$ is trained on the second half images of the batch, each containing about 15,000 images. Therefore, every feature of every labeled image from the ImageNet-1k training set is trained at a unique FNN node of the model diagram Fig.1. Thus, we can use notation $B_{j,s}$ to denote the batch, corresponding to label module $j$, $1\leq j\leq k$, and subset batch $s$, $1\leq s\leq r$. The partition of the training set applies to all features, $f_{i}$, $1\leq i\leq n$. As a result, in terms of feature $f_{i}$, the input image batch $B_{j,s}$ is transformed into the corresponding feature, which we denote as $B_{i,j,s}$. We can also denote the corresponding FNN by $N_{i,j,s}$, which is trained on the featured image batch $B_{i,j,s}$. That is, each FNN $N$ is independently trained on feature $f$ of its training batch $B$.
For all FNNs, $N_{i,j,s}$, the ReLU activation function is used for the hidden layer, the softmax is used for the output layer, and the cross-entropy loss function is used for the classification of FNN’s output. As for the FNN parameters in edge weights, we will use $W^{\ell}_{i,j,s}$ for $N_{i,j,s}$ with superscript $\ell=1,2$ for the first and second layers. The same notation convention, $b^{\ell}_{i,j,s}$, applies to the bias parameters of the FNN models. $L_{i,j,s}$ denotes the cross-entropy loss function of the FNN $N_{i,j,s}$. For a deeper FNN, $\ell=1,\dots,q$, for $q>2$.
By definition, to train the FNN model, $N_{i,j,s}$, is to find the global minimum of the loss function, $L_{i,j,s}$, at which every training image from $B_{i,j,s}$ is perfectly trained with the ground-truth label equal to FNN’s predicted label. We achieve this goal in two steps. In step one, we use a stochastic gradient descent (SGD) method to find a local minimum of the loss function. In step two, we use the gradient descent tunneling method ([5]) (GDT) to move from the SGD local minimum in the $\{W,b\}$ parameter space to a global minimum. In the end, we obtain two types of trained models: the SGD-trained proto-model, S_hxx_myy, represented by its local minimum $\{W_{s},b_{s}\}$, and the GDT-trained proto-model, T_hxx_myy, represented by its global minimum $\{W_{t},b_{t}\}$.
If the images, $B_{i,j,s}$, are self-consistent, namely, the labeling of images defines a function, the GDT method guarantees to find the global minimum, at which every image is correctly predicted by the FNN, $N_{i,j,s}$. If, on the other hand, the image labeling of the batch $B_{i,j,s}$ fails to be a function, then there exists at least one image that appears in $B_{i,j,s}$ at least two times with two different labels, i.e., a problem of double labeling. When this inconsistency occurs, the FNN is impossible to reach 100% accuracy. As a result, the GDT method will find a parameter point in $\{W,b\}$ at which the accuracy is below 100%. In other words, whenever the GDT-trained error rate is not zero, double labeling is suspected.
4. Evaluation
Once the proto-models are trained, their FNN parameters $\{W,b\}$ completely define the models. Two types of evaluation can be done: one is referred to as the training evaluation, and the other is referred to as the model evaluation. For the first type, it is straightforward: evaluate the performance of each FNN in terms of its accuracy rate on its assigned training data. In particular, whether the GDT-type rate achieves 100% accuracy or the training data contains contradictions. For the second type, all that matters is the proto-model defined by the parameters, regardless of how the parameters are obtained, for which the model’s architecture becomes important. Evaluation compares the predicted label by the proto-model of all FNN modules with the true label for every training image. In summary, training evaluation compares the true label only with the predicted label by the designated FNN module for the image, whereas model evaluation compares the true label with the predicted label from all the FNN modules by the majority-voting protocol.
In the ideal situation, if the entire training set of images is free of double labeling, then every image is perfectly trained for every feature. Thus, at the proto-model’s output node for model evaluation, there will be at least $n$ correct predictions, all having the same ground-truth label. This takes place at the same module number $j$ and the same subset number $s$ for all features. For any combination other than the intended FNN, $N_{i,j,s}$, for all $1\leq i\leq n$, the prediction can be any label from 0 to 999. By chance, these predictions can achieve a majority equal to the possible maximum, the number of features, $n$, referred to as the super-majority number. However, the larger the number of features $n$ is, the less likely a chancy super-majority becomes.
In fact, it is straightforward to see that if each feature is trained on only one module with $k=r=1$, then perfectly trained FNNs, $N_{i}:=N_{i,1,1}$, automatically lead to a perfectly trained model for all $1\leq i\leq n$, because majority voting will always output the true label, $n$ times, with the super-majority. If $k>1$ or $r>1$, then the aggregated model may not reach the 100% accuracy rate, because every training image is also subject to FNNs not trained on the image, which can result in a false label. In fact, when $r=1$ and $k>1$, the FNNs of modules, to which the image does not belong, always result in a false label, which can form a super-majority by chance. Therefore, the more features a model has, the higher its majority-ruled accuracy rate becomes, because it is less likely that false labels form a voting majority.
Figure 2. Confusion Wheel. The plot is for the proto-model S_h2_m1 and T_h2_m1, with module $m_{9}$, sub-module $n_{92}$, and feature $f_{4}={\rm RGg1}$, namely, FNN $N_{4,9,2}$. The 25 labels of the featured-module FNNs are represented alphabetically. Labeled output vectors of the FNNs are linearly transformed to corresponding spokes in the unit circle. A misclassified point lies outside its designated sector with the corresponding centering spoke. The linear transformation matrix consists of the spoke vectors on the unit circle for its column vectors.
To illustrate this analysis, we will use six featured models in the section below, each having a different feature combination, all derived from the proto-models. For Model-1, we use the first three features of Tab.1, i.e., the primary color channels of the images. Model-2 consists of the first six features of Tab.1, and the next of the first nine, and etc, up to Model-6, having all seventeen features from the table.
5. Results
Training Evaluation. In the GitHub repository [6], in the ImageNet-1k/Data_Transformation directory, the code, Run_Modularize_Data.m reorganizes the dataset into $k=40$ modules, and then partitions each module into $r=2$ groups of training images. For the training accuracy rate, the code, Run_Training_Evaluation_Model.m, evaluates the accuracy for the SGD-type and the GDT-type FNNs, $N_{i,j,s}$, with $1\leq i\leq 17,1\leq j\leq 40,1\leq s\leq 2$, on the modularized training images $B_{i,j,s}$, This is done by first transforming the original $64\times 64$ color images into the input vectors of the first layer of the FNN, and then obtaining the predicted labels of the FNN.
The transformation first down-samples the images to $32\times 32$, using the mean values of non-overlapping $2\times 2$ grid cells. Then it trims off one row from the top and bottom, and, respectively, one column from the left and right borders of the images. Featured images are obtained from the primary color channel images according to the combination weights from Tab.1. Each featured image is then scaled to cell values between [1,1][-1,1]. As a result, the final $30\times 30$ image matrix is changed to a 900-dimensional input vector for the FNN. The image transformation is performed by the function code, fun_transform_data_rgbfeatures.m. In the Data_Transformation folder, the code script, Run_Transform_Data.m, transforms the modularized dataset to the vectorized input for all training FNNs.
proto-model min mean median max 100rt stat.
T_h1_m1 98.965 99.923 99.975 100 196/1360, 12/80
S_h1_m1 98.866 99.815 99.876 99.981
T_h1_m2 98.736 99.919 99.969 100 196/1360, 12/80
S_h1_m2 98.723 99.831 99.888 99.994
T_h2_m1 98.884 99.898 99.956 100 154/1360, 10/80
S_h2_m1 98.745 99.783 99.856 99.987
Table 2. Training accuracy rates. All statistics are calculated against the total number of $N_{i,j,s}$, 1360. The last column is for the error-free FNNs, distributed over the total number of modules and their subsets, 80. For the T_h1_m1, for example, the number of error-free FNNs per module is 196/12 = 16.3333. That is, there are 11 modules of which all 17 features are trained error-free, and one of the 12 modules has 9 out 11 features that are trained error-free. The missing two feature FNNs can be made error-free by the GDT algorithm with stricter search settings.
Specifically, all SGD proto-models are trained with a threshold stop at 98%. Above the threshold, the GDT training follows right afterward to obtain the corresponding GDT proto-models. The SGD proto-model with two hidden layers, S_h2_m1, presented is of this throttling-down type. Figure 2 gives an illustration for these two types of proto-model conjoining training. It is chosen to show the dependency relationship between SGD and GDT trainings when the data batch $B_{i,j,s}$ is free of double labeling. In contrast, for the SGD proto-models with one hidden layer, S_h1_myy, $y=1,2$, after the GDT proto-models are trained, the threshold break for the SGD training is released, and we then train the SGD proto-models in a few more gradient descent iterations to a higher accuracy, allowing the training algorithm to realize its full potential. We use S_hxx_myy to denote this SGD training type, and $\{W^{\prime}_{s},b^{\prime}_{s}\}$ for their trained parameters.
Table 2 lists the performances for all types of training, in terms of the minimum, mean, median, and maximum accuracy rates, and the number of featured modules that reach 100% accuracy. For example, the result shows that of the 80 module partitions, 12 are free of doubling labeling. The minimum accuracy for the GDT training gives a lower estimate for the scope of the double-labeling problem. It also shows that the GDT proto-models always perform better than the SGD proto-models, with or without the SGD threshold break. The code, Run_Training_Analysis.m, is used to produce this table.
featured T_h1_m1 T_h1_m2 T_h2_m1
model Acc. Top-1 Acc. Top-1 Acc. Top-1
Model-1 86.848 99.998 86.047 99.999 73.065 100
Model-2 95.931 99.986 95.767 99.986 92.900 99.954
Model-3 97.382 99.982 97.275 99.980 95.979 99.866
Model-4 97.897 99.978 97.879 99.978 97.014 99.802
Model-5 98.133 99.979 98.126 99.970 97.487 99.755
Model-6 98.289 99.977 98.300 99.969 97.770 99.729
mean 95.747 99.983 95.566 99.981 92.369 99.851
S_h1_m1 S_h1_m2 S_h2_m1
Model-1 88.728 99.987 89.191 99.991 72.925 99.992
Model-2 96.282 99.757 96.371 99.796 91.570 99.715
Model-3 97.444 99.509 97.486 99.596 94.881 99.318
Model-4 97.905 99.309 97.953 99.429 96.068 98.991
Model-5 98.106 99.143 98.168 99.275 96.633 98.734
Model-6 98.247 99.008 98.299 99.166 96.990 98.569
mean 96.119 99.452 96.245 99.542 91.511 99.220
Table 3. Accuracy rates for featured models. Notice that Model-6 is identical to its proto-model because both contain all the training features.
Model Evaluation. As discussed above, when an image is fed to the model, the model must run it through all featured FNNs against the label classes 0 through 999 before making its prediction by the majority-vote protocol. For a single or a small batch of images, this is done by the code, Ru_Evaluation_Model_Individual.m, in the Model_Evaluation folder. But, for the entire dataset of 1.2 million plus images, this program is too slow. Instead, a parallelized code, Run_Evaluation_main_Model.m, is used to run on batches of the dataset to speed up the process. It outputs the predicted labels of the dataset for the feature-defined Model-1 through Model-6, introduced in the previous section. After the predictions are saved, we can run two more files. One file, Run_Evaluation_post_Model.m, finds the accuracy rates of all six models, and the Top-1 accuracy rate, which is defined to be the rate by which a correct prediction is made by the super majority of the votes, namely the first choice of the predictions. The other file, Run_Evaluation_Confusion_Matrix.m, builds the confusion matrices for the models, from which we can find the labels whose entire image classes are classified perfectly. All these tasks for all featured models, derived from the two types of proto-model can be accomplished by the code, Run_Model_Analysis.m in the ImageNet-k folder. The results are shown in Tab.3. It shows the following: The more features a model has, the more accurate it becomes, with a modest drop in the Top-1 rate. Models with a single hidden layer perform better than models with two hidden layers. The Top-1 rate for the GDT-trained featured models is always better than the SGD-trained featured models. For the conjointly trained models for which GDT training is directly based on SGD trained models, featured models of the former always perform better than models of the latter type, as shown by the h2-models. However, when this dependence is removed, i.e., the SGD proto-models are trained beyond the 98% threshold stoppage, their offspring featured models can outperform the GDT kind in model accuracy, but not in the Top-1 accuracy.
Errorless Trained Labels
featured T_h1_m1 T_h1_m2 T_h2_m1
model by batch all by batch all by batch all
Model-1 0.8 0 1.4 0 0 0
Model-2 59.7 0 46.9 0 4.6 0
Model-3 153.7 1 131.2 1 38.5 0
Model-4 220.9 7 201.8 5 90 0
Model-5 256.8 9 245.6 5 128.9 0
Model-6 285.9 11 278.9 8 168.3 1
Table 4. Confusion matrix statistics. Here is what the numbers mean. For example, for Model-6 of T_h1_m1, the ‘by batch’ number is the average number of labels that are perfectly classified over the 10 batches of the ImageNet-1k dataset, which is 285.9. In comparison, when aggregated over the 10 batches, the number drops to 11 labels, in the ‘all’ column. The corresponding labels are label-number 6, 293, 386, 420, 451, 467, 567, 640, 642, 644, and 708.
Table 4 summarizes some statistics of the confusion matrices. The code, Run_Confusion_Matrix_Analysis.m, generates all the numbers in the table. The main finding is that the more features a model has, the more perfectly classified labels the model is capable of gaining. It also shows that having one additional hidden layer degrades the outcome in all the evaluation categories analyzed in this paper. The same script can be used to find the statistics for all featured models of the SGD and GDT types.
6. Discussion
The continuous UAT ([9, 4]) holds a deep theoretical interest on its own. But, the discrete UAT is extremely important for applications because all practical problems can be treated as either discrete directly or approximated as discrete with the approximating error smaller than the UAT’s discriminating threshold. That is, all functions can be represented exactly by FNNs by the discrete UAT. The GDT method is the only known method capable of realizing this potential. Models that can be trained error-free are models that can truly improve on themselves.
The analysis and results from previous sections suggest that for datasets free of double labeling, the architecture with $k=r=1$ always gives rise to perfectly trained models. This can be the desired outcome if the dataset is not too big, such as the MNIST, and MedMNISTs from [20]. But for large datasets, such as the ImageNet-1k considered here, training a single FNN without modularization becomes prohibitive in terms of computer memory size and running time because the number of training images and the size of the output layer are both substantial. This is the primary reason why we adopted the modularized architecture for the model. In addition to the intended advantage of only requiring modest computing capabilities for training of modularized FNNs, model updating can be carried out by retraining the modules that need updating. For large datasets, we trade some loss of accuracy for efficiency by using modularization.
For our model, the modules are organized in the numerical order of the dataset’s labels. However, a more rational approach should be taken, similar to the idea of taxonomy. For example, we can group animals, plants, etc., in different modules. Another idea is to construct expert models, each for a distinct category of objects. The obvious advantage is efficiency because we do not need to check a known type of image against irrelevant types. Another obvious advantage is cost savings for retraining, for which only the relevant modules are updated. Datasets similar to ImageNet-1k, in our view, should be organized by categories.
For model architecture, our results suggest that having more hidden layers does not improve model performance. As a probable explanation, the nonlinear transformations of the additional layers may have erased some information essential for errorless classification. This finding seems consistent with the discrete version of UAT, which requires only one hidden layer for errorless classification. However, for datasets free of double labeling, GDT guarantees 100% accuracy regardless of the number of hidden layers.
We did not reach our goal of training the ImageNet-1k to perfection. The problem may have an explanation because we limited all the model parameters to 4-decimal-place accuracy. Although it is a possibility, it is unlikely because when our GDT algorithm runs at its maximum strength, the error-free trained modules are very uniform: either all features of a module are trained error-free, or no feature of a module can reach 100% accuracy, suggesting that some modules are free of double labeling and some others are not. In addition, the trained models with the given decimal place precision are robust with respect to the model parameters because all the results remain the same quantitatively when the parameters are truncated to the 3rd decimal place. As a result, we strongly believe that the ImageNet-1k may contain double-labeled images. However, it is not our goal to make a definitive determination.
Assuming that is the case, here is an estimate on the scope of the problem. From the minimum accuracy rates of the GDT proto-models in Tab.2, we can obtain an estimate for the maximum error rate around 0.0115, which translates to an estimated 14,000 images. Assuming only 10 percent of these images that is ultimately responsible for the double labeling problem, we are left with a set of 1,400 images. Averaged over 1000 labels, it suggests that there are 1 to 2 images per label on average that are double labeled, looking like the result of human curating error. An algorithm can be developed based on our GDT search algorithm to find the errant labels. This can be the subject of a future study.
If double labeling is detected in a dataset, we should be extra cautious when deploying a model trained on it. This is because the problem is bound to cause classification errors, which can cascade down to applications. For this reason, we did not pursue a model evaluation on the validation data since our search algorithm is not equipped to isolate the causes of the prediction error. The dataset’s probable curating errors can cause needless inaccuracy for the GDT-trained models, which can be amplified by the curating errors when the model is applied to the validation dataset. As a result, we used only the simpler kinds of features, consisting of linear and reversible combinations of the primary color channels, to test the GDT method on the dataset.
For image datasets free of double labeling, our GDT method can train models error-free, a vital goal in some applications, especially in the fields of Medicine and Healthcare, where repeating errors must not be tolerated. For problematic image datasets, our results suggest that it may not be worthwhile to add GDT training to SGD training, because the overall performance of our SGD models is on par with the GDT models, and GDT training is considerably more time-consuming than SGD training. We note that our SGD models with six features (Model-6), 113,798,880 parameters, and 566,880 neurons, have already outperformed all conventional SGD methods in the literature.
For practical applications and image datasets free of double labeling, we envision the inclusion of more sophisticated features, such as various convolutions of, or singular value decompositions of images. Such feature inclusions are for future studies to quantify the predictive capability of trained models in terms of the size of neighborhoods of training data in which models’ predictions remain true. Such studies should have a more concise explanation for models’ performance on validation and test data, as well as the true scope of models’ capabilities and limitations. In addition, only for consistent datasets can we pursue the smallest FNN model that can be trained fully. For the MINST handwritten digits, we only need one hidden layer of 20 neurons and 15,910 parameters to achieve that goal ([5]). We believe that the size of our model for ImageNet-1k can be reduced by one order of magnitude for the same level of performance, which we did not pursue in this study.
The tree-like structure of our model is similar to taxonomy. We believe that datasets, such as species taxonomy, the catalog of the Library of Congress, or any database requiring precision in bookkeeping, can be represented by AI models like ours. Most importantly, if such a model makes a mistake, the error-free training method can ensure that an updated module will not make the same mistake. Errorless training can be beneficial for human endeavors that require predictive precision AI.
Because our model is structured in parallel FNNs with minimal hidden layers and modest batch sizes, training is limited only by the number of CPU cores. In the ideal setting where each FNN is assigned to a core, an SGD proto-model can be trained in 1 hour, and a GDT proto-model can be trained in 3 hours, which is the same length of time to update any FNN. As a demonstration, no GPUs were used to train our model.
Acknowledgment: All large-scale computations for the paper were carried out in the Holland Computing Center of the University of Nebraska System. Thanks to Lucas Wurtz for his enthusiasm to get the project started.
Data Availability: All model weights and evaluation codes can be found in the GitHub repository, [6]. For questions about the training methods, send emails to the first author.